# Past, Present, and Future of Citation Practices in HCI

### v1

@@ 4.3.3. Pre-print articles are increasingly being cited by CHI authors
The landscape of academic publishing is changing. In this section, we examine how CHI authors cite pre-print articles. Pre-print articles are versions of scholarly articles that precede publication in peer-reviewed journals and conference venues. While traditionally shunned upon due to not being formally reviewed, it has become common place in some research fields, such as Machine Learning, to upload and also cite pre-print articles. Pre-print archives offer a convenient way for authors to claim priority capital, establishing the authors’ precedence in scientific discoveries or novel ideas (Newman, 2009). One of the most popular pre-print archives is arXiv, launched by Cornell University in 1991. In recent years, an exponentially growing number of articles are being submitted to this pre-print archive. The high number of available pre-prints and the success of pre-prints in certain fields, such as Machine Learning, may incentivize CHI authors to include more citations to pre-prints in their CHI articles. To investigate this in more depth, we scraped the references from the ACM–DL for each article in the CHI proceedings. We then identified the number of citations to arXiv by counting the number of times “arxiv” is mentioned in the articles’ reference section. For counting the number of occurrences of arXiv in the references section, we used a case-insensitive regular expression that accounts for hyphenation within the word arXiv due to line breaks.
[…]
Number of CHI articles citing arXiv pre-print since 2006.
Figure 9. Citations to pre-print arXiv papers (a) and percentage of CHI articles citing arXiv pre-prints $n$ times (b) in the CHI Proceedings.
@@
Even though arXiv was made available to the World Wide Web in 1993, we find that none of the CHI articles before the year 2006 include citations to arXiv pre-print articles. 9(a) depicts the number of citations to arXiv pre-prints in CHI articles since 2006. We note there is a change in this time series after 2015. While the mean number of citations to arXiv pre-print articles has only slightly increased since 2016 (from 0.03 references in 2015 to 4.24 references in 2024), citations to pre-print articles have become overall more accepted among some CHI authors (9(a)). One article in CHI ’24 included 89 references to arXiv. 9(b) depicts the number of times $n$ that an article cites arXiv articles. From this figure, one can note an increase in the number of articles that cite a high amount of arXiv pre-prints. In recent proceedings, over a third of the CHI papers include more than one citation to arXiv. Given that this change coincides with the year 2016, citations to arXiv articles are one contributing factor to the increase in references per CHI article.
4.3.4. Data-driven research is on the rise
With the rise of big data and data-driven research methodologies, CHI papers may have started referencing more data sources, scientific software and tools, code repositories, and prior data-heavy studies. We investigate this trend by counting the number of references to five common open data and source code repositories:
  • osf.io, a cloud-based management solution for open access science, often used to share datasets and other data related to HCI studies,
As above, we limit our investigation to the articles’ reference sections which we scraped from the ACM–DL. [^3] We ignore citations in footnotes within the article’s full text to make the results comparable to the other findings in this article.
[…]
Mean number of citations to data and code repositories per CHI article.
[…]
Relative number of CHI articles citing data and code repositories at least once.
Figure 10. Mean number of citations to code and data repositories (a) and relative number of CHI articles citing at least one repository (b) in the CHI proceedings from 2008 to 2024.
[…]
Mean number of citations to data and code repositories (a) and relative number of CHI articles citing at least one repository (b) in the CHI proceedings from 2008 to 2024.
@@ 4.3.4. Data-driven research is on the rise
10(a) plots the mean number of references to code and data respitories in CHI articles. While the average number of citations to such repositories per CHI article is still low, we identify a clear increase in the number of citations to code and data repositories occurring after the year 2015. Before 2015, almost no authors formally cited these repositories, whereas today, such citations have become more common. In the proceedings of CHI 2024, 147 articles (13.9%) cite at least one data and code repository (see 10(b)). GitHub was cited most often with 115 articles (10.8%) in the CHI 2024 proceedings including at least one reference to this repository. This empirical trend in the citation practices reflects the broader trend towards applying data-driven methodologies and machine learning in HCI research.
@@ 5. Discussion
This article presented meta-research on the field of HCI, providing evidence of clear quantitative and qualitative changes in the citation practices of the CHI community co-occurring since the year 2016. Specifically, we have described five co-occurring changes:
@@
    In the 2016 Call for Papers, references were no longer included in the page limit of CHI articles. This editorial decision launched the CHI community on an expansive path, characterized by a year-by-year expansion of the mean number of references included in CHI articles. One could argue that the observed trend in Figure 1 is simply an effect of the overall growth of publications in the HCI field. However, while growth may certainly be a contributing factor, the growth of the CHI conference proceedings precedes the observed trend by 10 years, as evident in Figure 2. Therefore, the increase in the mean number of references per CHI article is likely not a mere side-effect of CHI proceedings growth. Rather, it is the top-down policy change that took place in the year 2016 that opened the floodgates, allowing authors to cite more and more works in their articles.
@@
    An increase in collaboration and the number of references per author (Section 4.3.1):
    It is becoming more common for authors in HCI to collaborate and the mean number of authors per CHI article is expanding. An increase in the number of co-authors means that the set of authors may potentially contain authors from different disciplines. This implies that authors potentially draw from a broader range of interdisciplinary sources, hence leading to more citations to other articles. However, our analysis found that the increase in the mean number of authors is not proportionate to the increase in the mean number of references, as evident in Figure 7. The growth in the mean number of references exceeds the growth in authors.
@@
    The literature is growing at a rapid rate, and we found that CHI authors increasingly conduct systematic literature reviews of the expanding corpus of scholarly literature. The COVID-19 pandemic years saw an acceleration of this trend, denoted by an increase in the relative number of CHI articles conducting literature reviews. This can be explained by authors not being able to conduct user studies, a common method in HCI, during the pandemic. However, as depicted at the top in Figure 8, the growing trend toward literature reviews already started earlier. The editorial decision to lift the page restrictions greatly contributed to the growing popularity of literature reviews at CHI. Literature reviews typically contain an above average number of references, which contributes to the increase of the mean number of references included in CHI articles.
@@
    An increase in the number of citations to arXiv (Section 4.3.3):
    Entire fields of science rely on arXiv as a means for article dissemination. The field of HCI is not immune to this trend, and it is becoming increasingly more common for CHI authors to cite unrefereed articles. Before 2015, citations to arXiv were shunned upon by CHI authors. From 2016 onward, citations to arXiv were no longer competing for space with citations to peer-reviewed articles. The editorial decision of lifting page restrictions in CHI articles contributes to this shift in citation behavior. Today, many CHI authors cite arXiv pre-print articles. This can lead to adverse problems, such as the pre-print being cited instead of the published article.
@@
    Before 2015, there were virtually no formal citations to code or data repositories to be found in the references sections of CHI articles. From our analysis, it can be concluded that lifting the page restrictions animated authors to include more citations to such repositories in their CHI articles. From an economic lens, there no longer is a marginal cost to including another citation to arXiv or code and data repositories and there is no trade-off cost between citations.
@@
Together, these co-occurring trends shape the citation practices of the CHI community in the shifting landscape of HCI. These trends reflect an interesting change of pace in academic working in the CHI community starting with CHI ’16. The trends are also a reflection of broader changes that affect not only the field of HCI, but academia in general.
@@
What is striking about the observed trend in Figure 1 is that it follows a relatively stable linear growth pattern. This near-linear growth takes place in the presence of exponential growth in academia, such as the exponentially growing number of scholarly articles. Future work could investigate why the observed year-by-year growth in references is linear. Perhaps it could be explained with a simple psychology model of herd mentality. Each year, members of the CHI community observe — whether consciously or sub-consciously — an increase in the mean number of references per CHI article. In other fields of science, evidence has been found that longer articles are cited more (Kousha and Thelwall, 2022). A simple strategy of “the more, the better” could be motivating authors in CHI to include more references in their articles to satisfy the seemingly growing expectations of the peer reviewers. In the following section, we discuss why the growing number of references is a critical concern that demands attention.
@@ 5.1. Implications for Authors, Reviewers, and the Broader Academic Community
An article with many references can help novices and those unfamiliar with a research area to learn and become familiar with the research area. However, we argue the observed trend (i.e., excessive citations as evident in the rising number of references in CHI articles) is a negative development for authors, peer reviewers, and the entire field of HCI, for a number of different reasons.
@@
First, we argue that 130 references per article is too many to thoroughly screen during the peer review process. A high number of references increases the cognitive load on peer reviewers, contributing to reviewer fatigue. This only adds to the already high amount of peer review fatigue, with conferences struggling to recruit high-quality reviewers. Peer review is the backbone of academic rigor, and if reviewers only give a cursory glance at the reference section, the integrity of the peer review process is compromised.
@@
Second, the number of references included in CHI articles serves as a signal to the CHI community. Peer pressure and herd mentality suggest that if the average number of references is high, authors feel compelled to include a similar number of references in their articles. Deviations from this norm are likely to be viewed negatively by peer reviewers, creating a spiraling pattern of year-by-year increasing academic pressure and competition. This spiraling pattern contributes to author dissatisfaction and fatigue.
@@
Third, despite the shift away from awarding best paper awards and honorable mentions to high-reference articles, the overall rise in the number of references still sends the wrong message to junior members of the CHI community. Today, PhD students in HCI are faced with a seemingly insurmountable and overwhelming amount of literature work, which distracts them from focusing on developing their own research skills and expertise.
@@
Fourth, the inflation of citations diminishes the value of each citation. A citation today does not carry the same weight it did in the past. This is similar to how early search engine optimization efforts devalued the significance of web links. The increasing number of citations reduces their impact and can lead to an overload of less meaningful references.
@@
Last, the high number of references makes it easier for bad actors to commit fraud. In an environment of hypercompetition and “perverse” incentives (Edwards and Roy, 2017), academic fraud is surprisingly common (Tijdink et al., 2014). With no clear trade-off cost between citations in an article, the reference section can be expanded arbitrarily, facilitating citation manipulation. This can include self-citation (Bartneck and Kokkelmans, 2011), author collusion in citation cartels (Kojaku et al., 2021; Jolly et al., 2020), and even the purchasing of citations (Ibrahim et al., 2024). The high number of references that, likely, undergo only cursory peer review, makes it harder to detect such unethical and fraudulent research practices.

### v7

@@ 4.2.4. Pre-print articles are increasingly being cited by CHI authors
Even though arXiv was made available to the World Wide Web in 1993, we find that none of the CHI articles before the year 2006 included formal citations to arXiv pre-print articles. Figure 11 depicts the number of citations to arXiv pre-prints in CHI articles since 2006. We note there is a change in this time series after 2015. While the mean number of citations to arXiv pre-print articles has only slightly increased since 2016 (from 0.03 references in 2015 to 4.24 references in 2024), citations to pre-print articles have overall become more accepted among some CHI authors. One article in CHI ’24 included 55 references to arXiv. Figure 11 depicts the relative number of articles citing arXiv pre-prints $n$ times, with $n$ split into buckets (1 citation, 2–10 citations and more than 10 citations). From this figure, one can note an increase in the number of articles that cite a high amount of arXiv pre-prints (¿10 citations). In recent proceedings, over a third of the CHI papers include more than one but less than 10 citation to arXiv. Given that this change coincides with the year 2016, citations to arXiv articles are one potential factor contributing to the increase in references per CHI article.
4.2.5. Data-driven research is on the rise at the CHI Conference
@@
Figure 13 plots the mean number of references to code and data repositories in CHI articles. While the number of formal citations to such repositories per CHI article is still low, we identify a clear increase in the number of citations to code and data repositories occurring after the year 2015. Before 2015, almost no authors formally cited these repositories, whereas today, such citations have become more common. In the proceedings of CHI 2024, 147 articles (13.9%) cite at least one data and code repository (see Figure 13). Zenodo and Huggingface were cited too infrequently to include in figures 13 and 13. GitHub was cited most often with 115 articles (10.8%) in the CHI 2024 Proceedings including at least one reference to this repository. This empirical trend in the citation practices reflects the broader trend toward applying data-driven methodologies and machine learning in HCI research.
4.2.6. CHI authors increasingly cite questionable publishers.
We identified 3,298 citations from CHI articles to potential predatory journals and publishers in the CHI Proceedings (1981 – 2024). There is a marked increase in citations to potential predatory journals and questionable publishers since 2016, both in absolute and relative numbers (see Figure 14a and Figure 14b, respectively).
Citations to Frontiers Media – the publisher which successfully pressured Jeffrey Beall to shut down his list – account for the vast majority of these citations. Splitting Frontiers from the rest of potential predatory publications, we can see that citations from CHI articles to journals published by Frontiers Media have monotonously increased since 2015, both in absolute and relative terms (see Figure 14e and Figure 14f). In 2024, every CHI paper includes on average about 0.8 references to a Frontiers journal (cf. Figure 14f). The journals from Frontiers Media cited most often in the CHI Proceedings are Frontiers in Psychology ($n=873$, 39.72% of all citations to Frontiers Media), Frontiers in Human Neuroscience ($n=248$, 11.28%), Frontiers in Robotics and AI ($n=213$, 9.69%), Frontiers in Virtual Reality ($n=130$, 5.91%), Frontiers in Neuroscience ($n=86$, 3.91%), Frontiers in Psychiatry ($n=68$, 3.09%), Frontiers in Public Health ($n=67$, 3.05%), Frontiers in Computer Science ($n=55$, 2.5%), Frontiers in ICT ($n=37$, 1.68%), Frontiers in Neurology ($n=32$, 1.46%), Frontiers in Artificial Intelligence ($n=31$, 1.41%), and Frontiers in Behavioral Neuroscience ($n=26$, 1.18%). This is a testament that the CHI community has embraced Frontiers Media as a legitimate publisher.
Table 3. The results of an event study indicate a significant change in the mean number of references (dependent variable) per CHI article before (2007–2015) and after (2016–2024) the decision to lift page restrictions at the CHI Conference.
Statistic Value
R2 0.999 dfregression 9
Adj. R2 0.997 dfresiduals 8
F-test $F=687.6$, $p<1e^{-9}$
Parameter Coefficient std err t p CI (95%)
Coefficient $\beta_{0}$ (intercept) 34.9467 2.498 13.992 0.000* [29.187, 40.706]
Coefficient $\beta_{1}$ (indicator) 6.5244 1.771 3.683 0.006* [2.439, 10.609]
Coefficient $\beta_{2}$ (AwardRatio) 1.9100 5.631 0.339 0.743 [-11.075, 14.895]
Coefficient $\beta_{3}$ (Authors) -9.2715 5.969 -1.553 0.159 [-23.035, 4.492]
Coefficient $\beta_{4}$ (Arxiv) -3.6121 1.651 -2.188 0.060 [-7.419, 0.195]
Coefficient $\beta_{5}$ (Repos) -2.7982 29.674 -0.094 0.927 [-71.227, 65.631]
Coefficient $\beta_{6}$ (Reviews) 12.8483 48.702 0.264 0.799 [-99.459, 125.156]
Coefficient $\beta_{7}$ (Predators) 38.7569 10.727 3.613 0.007* [14.021, 63.493]
Coefficient $\beta_{8}$ (Interaction) 3.3381 0.999 3.341 0.010* [1.034, 5.642]
Coefficient $\beta_{t}$ (Time) 1.8930 0.535 3.537 0.008* [0.659, 3.127]
Besides Frontiers Media, the remaining matches are mostly one-off citations to potential predatory journals and publishers. While there is an increase in the absolute number of citations to these potential predatory journals (other than Frontiers; see Figure 14c), the relative number of citations remained stable and limited in the recent decade (Figure 14d). Among these citations, IOS Press ($n=418$) and IGI Global ($n=195$) are occurring most often, followed by IntechOpen ($n=36$), International Journal of Computer Applications ($n=28$), Scientific World ($n=15$), Journal of Physical Therapy Science ($n=13$), Cambridge Scholars Publishing ($n=9$), Global Media Journal ($n=8$), and Research and Reviews ($n=8$).
4.3. Factors Explaining the Rise in the Mean Number of References (RQ3)
Our event study provides insights into the significance of the change observed in this work and the factors influencing the mean number of references in CHI articles across the study period (see Table 3). The indicator variable for the post-event period ($\beta_{1}$) shows a statistically significant effect ($p=0.006$), suggesting a notable shift in the mean number of references after the policy change. The interaction term ($\beta_{8}$) is also significant, indicating a significant change of slope post-event.
Among the other variables, the awards given at the CHI Conference to high-reference articles ($\beta_{2})$ had no statistically significant effect on the mean number of references in articles. As visually demonstrated in Figure 7, the mean number of authors ($\beta_{3}$) also had no significant effect on the mean number of references. Among the other coefficients, citations to potential predatory publishers ($\beta_{7}$) is statistically significant ($p=0.007$), indicating that changes in predatory citation patterns may play a meaningful role in shaping citation practices among CHI authors. This can be largely attributed to the publisher Frontiers Media becoming a mainstay in the references of CHI articles since 2015. Conversely, the coefficients for citations to arXiv ($\beta_{4}$) and repositories ($\beta_{5}$), as well as literature reviews ($\beta_{6}$) are not statistically significant, suggesting these factors may have a limited direct impact on the mean number of references in CHI articles. The time trend coefficient ($\beta_{t}$) is statistically significant, pointing to other potential underlying effects explaining the trend in the mean number of references per article. These findings highlight the complexity of citation dynamics and the multifaceted factors influencing scholarly referencing practices. As mentioned in Section 4.2, this investigation is not exhaustive and there could be multiple other potential relationships that warrant exploration in future work.
@@ 5. Discussion
This article presented meta-research in the field of HCI, providing clear descriptive evidence of quantitative and qualitative changes in the citation practices of the CHI community co-occurring since the year 2016. Specifically, we have described seven co-occurring changes in citation practices:
@@
    In the 2016 Call for Papers, references were no longer included in the page limit of CHI articles. This editorial decision launched the CHI community on an expansive path, characterized by a year-by-year increase in the mean number of references included in CHI articles. Our event study specification provides empirical evidence of a significant change in citation behavior, with the CHI community departing from established patterns in between the years 2015 and 2016. The top-down policy change opened the floodgates, allowing authors to cite more and more works in their articles.
@@
    Systematic bias in awarding articles post-policy change (Section 4.2.1):
    A systematic bias in awarding articles was observed. In the years following the policy change (with exception of 2016), articles with above-average number of references were systematically presented with awards. However, the event study found this systematic bias in awarding articles not to be a statistically significant factor in predicting the mean number of references in CHI articles.
  3. (3)
    An increase in collaboration and the mean number of references per author (Section 4.2.2):
    It is becoming more common for authors in HCI to collaborate and the mean number of authors per CHI article is expanding. An increase in the number of co-authors means that the set of authors may potentially contain authors from different disciplines. This could imply that authors potentially draw from a broader range of interdisciplinary sources, hence leading to more citations to articles from different fields. However, both our visual and statistical analysis found no evidence of growth in authorship being a significant factor contributing to the growth in the mean number of references in CHI articles, as also evident in Figure 7. The growth in the mean number of references exceeds the growth in authors (which has been quite linear). Author growth alone, therefore, does not serve as explanation for the increase in the mean number of references per CHI article.
@@
    The corpus of scholarly literature is expanding at a rapid rate, and we found that CHI authors increasingly conduct both systematic and unsystematic literature reviews. The editorial decision to lift the page restrictions in 2016 greatly contributed to the growing popularity of literature reviews at CHI. The years during the COVID-19 pandemic saw an acceleration of this trend, denoted by an increase in the relative number of CHI articles conducting literature reviews. This can be explained by authors not being able to conduct in-person user studies, a common method in HCI, during the pandemic. Literature reviews typically contain an above average number of references, which may contribute to the increase in the mean number of references included in CHI articles. However, as depicted in Figure 9, the trend toward literature reviews already started before the pandemic. The event study found that literature reviews do not have a statistically significant influence on the mean number of references. This can be explained by literature reviews, while growing in popularity, still being relatively rare in the CHI Proceedings.
@@
    An increase in the number of citations to arXiv pre-print articles (Section 4.2.4):
    Entire fields of science rely on arXiv as a means for article dissemination. The field of HCI is not immune to this trend, and it is becoming increasingly more common for CHI authors to cite unrefereed pre-prints. Before 2015, citations to arXiv were shunned upon by CHI authors. From 2016 onward, citations to arXiv pre-prints were no longer competing for space with citations to peer-reviewed articles. The decision to lift the page restrictions at CHI paved the way for this shift in citation behavior. Today, many CHI authors cite arXiv pre-print articles. This can also lead to adverse problems, such as pre-prints being cited instead of the published articles. While the trend of citing arXiv pre-prints was not statistically significant in the event study, the mean number of citations to pre-prints approaches significance ($p=0.06$), hinting at a potential relationship that could develop significance in the future.
@@
    Before 2015, there were virtually no formal citations to code or data repositories to be found in the reference sections of CHI articles. From our analysis, it can be concluded that lifting the page restrictions motivated authors to include more citations to such repositories in their CHI articles. From an economic lens, without a page limit, there no longer is a marginal cost to including formal citation to code and data repositories, and authors no longer have to trade-off between citations. The event study found this trend not to be statistically significant, which could be explained with the relative sparsity of citations to code and data repositories in the CHI Proceedings.
  7. (7)
    An increase in the number of citations to predatory journals and questionable publishers (Section 4.2.6):
    Without a trade-off cost between citations, authors are increasingly including citations to potential predatory journals and questionable publishers in their CHI articles. The long-tailed distribution of citations to publishers on Beall’s List in Figure 14 highlights that the CHI community, overall, has not lost its sense for quality. The relative number of citations to smaller potential predatory journals and publishers is negligible compared to the overall number of citations in each proceedings year. However, the picture is different with larger questionable publishers (Frontiers Media, IOS Press, IGI Global, and IntechOpen) which are increasingly being cited by CHI authors. Perhaps this is yet another expression of the CHI community’s obsession with quantity.
@@
Together, these co-occurring trends shape the citation practices of the CHI community in the shifting landscape of HCI. The trends reflect an interesting gradual change of pace in academic working in the CHI community starting with CHI ’16. The trends are also a reflection of broader changes that affect not only the field of HCI, but academia in general (Edwards and Roy, 2017).
Like a frog slowly boiling in water, the CHI community is heating up under the relentless growth of references, risking a culture where quantity overshadows meaningful discourse and genuine innovation. A common proverb states: “things need to get a lot worse before they get better”. One has to wonder: At what point will the number of references per CHI article become unbearable for authors and peer reviewers – at 100 references, 130 references, 200 references? The pre-2016 limit on the number of pages in CHI articles constituted a “Schelling fence” (Alexander, 2012) – a hard limit on the number of pages that could be included in CHI articles (including references). Authors could not cross this limit without risking a desk-reject of their paper. Thus, the space for references and content had to be weighed against each other, and each additional reference reduced the space available for the article’s content. Without these clear boundaries and trade-offs, the CHI community is on a slippery slope (Alexander, 2012), a gradual erosion – or evolution, if one prefers to look at it this way – of norms and principles. The expansive path incentivizes authors to, year-by-year, value quantity over quality and include more references in their articles.
One could argue that the observed trend in Figure 1 is simply an effect of the overall growth of publications in the HCI field. However, while overall publication growth may certainly be a contributing factor, the growth of the CHI Proceedings cannot fully explain the observed growth in the mean number of references for two reasons. First, the growth of the CHI Conference Proceedings precedes the observed trend by 10 years, as discussed in Section 2.1. Second, the CHI Proceedings deviated from linear growth during the COVID-19 pandemic, while the mean number of references in CHI papers still increased steadily and fairly linearly. Therefore, the increase in the mean number of references per CHI article is likely not a mere side-effect of the proceedings growth at CHI.
@@
What is striking about the observed trend in Figure 4 is that it follows a relatively stable linear growth pattern. This near-linear growth takes place in the presence of exponential growth in academia, such as the exponentially growing number of scholarly articles. Future work could investigate why the observed year-by-year growth in references is linear.
This near-linear growth could, perhaps, be explained with the simple psychological models of herd mentality and an availability heuristic. Each year, members of the CHI community observe — whether consciously or sub-consciously — an increase in the mean number of references per CHI article. In other fields of science, evidence has been found that longer articles are cited more (Kousha and Thelwall, 2022). Authors may, thus, decide to also include more references in their articles. Another explanation are the rising expectations of peer reviewers. The submission guide at CHI 2016 explicitly mentioned that “lack of references to prior work is a frequent cause for complaint – and low rating – by reviewers” (ACM SIGCHI, 2016). It is not uncommon for peer reviewers to demand inclusion of a long list of references in their reviews. Authors often interpret these references not as suggestions but as mandatory for paper acceptance, and subsequently incorporate them into their paper. A simple strategy of “the more, the better” could be motivating authors in CHI to – preemptively – include more references in their articles to satisfy the growing expectations of peer reviewers.
Likely, though, it is a combination of several different co-occurring factors such as the ones presented in this article (i.e., the increase in systematic reviews in HCI, growing diversity of topics and co-authorship, growing expectations among peer reviewers, growing number of publications, and pressure to publish) that contributes to the increase in the mean number of references in CHI articles. We leave a detailed qualitative investigation of the motivation of authors to include more references in their papers to future work (for instance with interview studies, workshops, or a Delphi study). In the following section, we discuss why the growing number of references in articles is a critical concern that demands attention.
@@ 5.1. Implications for Authors, Reviewers, and the Broader Academic Community
References in scholarly articles serve an important function: they help novices in the field to learn and become familiar with the research area. However, we argue the observed trend (i.e., a culture of excessive citation, as evident in the rising number of references in CHI articles) is a negative development for authors, peer reviewers, the CHI community, and the entire field of HCI, for a number of different reasons.
@@
First, we argue 130 references per article are too many references to thoroughly screen during the peer review process. A high number of references increases the cognitive load on peer reviewers, contributing to reviewer fatigue. This only adds to the already high amount of peer review fatigue, with conferences struggling to recruit high-quality reviewers. Peer review is the backbone of academic rigor, and if reviewers only give a cursory glance at the reference section, the integrity of the peer review process is compromised.
@@
Second, the number of references included in CHI articles serves as a signal to the CHI community. Peer pressure and herd mentality suggest that if the average number of references is high, authors feel compelled to include a similar number of references in their articles. Deviations from this norm are likely to be penalized by peer reviewers, creating a spiraling pattern of year-by-year increasing academic pressure and competition. This spiraling pattern contributes to both author and peer reviewer dissatisfaction and fatigue.
@@
Third, despite the shift away from awarding best paper awards and honorable mentions to high-reference articles, the overall rise in the mean number of references per article still sends the wrong message to junior members of the CHI community. The literature work required to publish an article at CHI poses a significant and undue burden on young scholars. Today, PhD students in HCI are faced with a seemingly insurmountable and overwhelming amount of literature, which distracts the students from developing their research skills and expertise.
@@
Fourth, excessive citation contributes to inflation of the value of each citation. Some authors may use “citation stuffing” to prop up the perceived value of their articles. Due to this carelessness, a citation today does not carry the same weight it did in the past. This is similar to how early search engine optimization efforts devalued the significance of web links. The increasing number of citations reduces their impact and contributes to a hemorrhage of less meaningful references.
@@
Last, the high number of references makes it easier for bad actors to commit misconduct and fraud. In an environment of hypercompetition and “perverse” incentives (Edwards and Roy, 2017), academic fraud is surprisingly common (Tijdink et al., 2014). With no clear trade-off cost between citations in an article, the reference section can be expanded arbitrarily, facilitating citation manipulation. This can include self-citation (Bartneck and Kokkelmans, 2011; Labbé, 2010), author collusion in citation cartels (Kojaku et al., 2021; Jolly et al., 2020), and even the purchasing of citations (Ibrahim et al., 2024; Else and Van Noorden, 2021). The high number of references that, likely, undergo only cursory peer review makes it harder to detect such unethical and fraudulent research practices.